OpenStudy (anonymous):

Algebraic expressions 1) Non-negative values a and b satasfyed a+b=0. Find the values of a and b. 2) Factor: (x^2-2x)^2-2(x^2-2x)-3

OpenStudy (accessdenied):

1) a + b = 0 If we solve this for one value a = -b If we make one positive, the other \(has\) to be negative. So I believe we only have one possible solution. Can you see it? 2) We should treat this as if w = x^2 - 2x: w^2 - 2w - 3 . Are you able to factor this quadratic?

OpenStudy (accessdenied):

And for #1 we should distinguish between non-negative and positive. They're not quite synonymous, because there is one number between positives and negatives that is non-negative but not positive.

OpenStudy (anonymous):

I dont really know what you mean by one number thats not negitive and not positive, it can only be a decimal right, but it also can be a decimal because itwont equal to 0

OpenStudy (accessdenied):

0 itself is neither positive nor negative. Even though it doesn't have a negative sign, it has no magnitude in the positive direction either. It is in between positive and negatives.

OpenStudy (anonymous):

Can't^

OpenStudy (accessdenied):

If you think about the positive numbers with negative numbers as their inverses for addition, and, a + -a = 0 For 0, what is its inverse? 0 + ? = 0 Well, 0. 0 + 0 = 0, 0 - 0 = 0 0 is its own inverse.

OpenStudy (accessdenied):

But it wouldn't make sense to call 0 both positive and negative, the groups are different. So 0 is neither positive nor negative. Does that make sense?

OpenStudy (accessdenied):

But then one more step ahead, non-negative means "not negative" the positives are not negative, that is true. zero, being neither positive nor negative, is ALSO not negative! :)